6.2 The e conference department of the MerryWeather hotel in Inverness, Scotland has fixed costs of 360 dollar per day. A local university is seeking a quote in connection with a conference it is planning to hold next year. The university would like the hotel to provide the morning coffee, lunch, afternoon tea and to prepare conference materials to be distributed to all all conference attendees. The hotel's cost of providing food and drink during the day is 7 dollars per attendee. In addition, preparing the conference materials would cost the hotel 6 dollars per attendee. the university has estimate that the conference will be attended by between 80 and 120 people

A— what is total average cost per attendee if the conference has an attendance of 120 people.


B— if 120 people attend the conference and the hotel wants to earn a profit of 20% on revenue ,what price per person should be charged.



C--There are several other local hotels that provide university, which holds several conferences per year, has always used a venue provided by MerryWeather's main competitor. MerryWeather's conference manager is exceedingly keen to develop a working relationship with the University and has received permission from the General Manager to provide a very aggressively priced quote to the University. To help in his deliberations on this matter, the conference manager has approached MerryWeather's accounting department and asked them to determine what lowest price that can be charged per attendee without the conference adversely affecting this year's hotel profit. If you were an accountant at MerryWeather, what advice would you give to the conference manager?
Answer #1

A)

Cost of food and drinks per attendee = $7

Cost of material per attendee = $6

Hence, variable cost per attendee = Cost of food and drinks per attendee + Cost of material per attendee

= 7 + 6

= $13

Hence, Variable cost of 120 people = 120 x 13

= $1,560

Fixed cost per day = $360

Hence, total cost of 120 attendees = Variable cost + Fixed cost

= 1,560 + 360

= $1,920

Total average cost per attendee = Total cost/Number of attendees

= 1,920/120

= $16

B)

Profit = 20% on revenue

Let the revenue be $Y

Profit = Y x 20/100

= 0.2Y

Revenue = Total cost + Profit

Y = 1,920 + 0.2Y

0.8 Y= 1,920

Y = 1,920/0.8

= $2,400

Hence, revenue = $2,400

Price per person = Total revenue/Number of persons

= 2,400/120

= $20

C)

Lowest price that can be charged per attendee without the conference adversely affecting this year's hotel profit will be equal to the variable cost per attendee.

Hence, lowest price per attendee = $13

Fixed cost of $360 per day is immaterial in deciding the lowest price to be charged since it will not be increased due to thus particular conference.

Kindly comment if you need further assistance. Thanks